# AgentLeak: A Full-Stack Benchmark for Privacy Leakage in Multi-Agent LLM Systems

### v1

@@ VI-E Results
Each of the 4,979 traces contains both a single-agent execution (C1 only) and a multi-agent execution (C1+C2+C5) of the same scenario on the same model. This paired design enables direct comparison. Proportion of traces in which at least one sensitive field leaked through a given channel is used in Tables VI, VIII, XII. For high-level comparison, we use the arithmetic mean of C2 and C5 (57.8%) as a representative internal rate. The combined OR rate ($C2\lor C5$) is 68.8%.
@@
TABLE VII: Defense Effectiveness by Channel Type (A0 Benign Scenarios Only). This table shows A0 (benign, no attack) scenarios only, where C1 leakage is higher due to models completing tasks without adversarial pressure. Under adversarial conditions (A1/A2), internal channels dominate as shown in Table VI. C1 Eff. = (Vanilla C1 Rate - Defense C1 Rate) / Vanilla C1 Rate $\times$ 100%.
@@
TABLE VIII: Architecture Comparison: Single-Agent vs Multi-Agent (n=4,979 paired traces). Statistical test: McNemar’s test for paired proportions. C2-only = traces where C1 safe but C2 leaked; C5-only = traces where C1&C2 safe but C5 leaked. C5-only refers to the incremental leak rate; the +0.0% indicates that in our dataset, every memory leak (46.7%) co-occurred with an inter-agent (C2) or output (C1) leak.
@@
Why the increase? Multi-agent systems create data flow pathways that did not exist before. Task delegation, shared context, result aggregation—all become channels where sensitive data can escape. The pattern is monotonic: add an agent, add a pathway. Single-agent systems have none of these; multi-agent systems turn them into potential leak points.
@@
What stands out is that C2 beats C1 for every single model we tested: GPT-4o (76.8% vs 17.2%), GPT-4o-mini (75.3% vs 41.2%), Mistral Large (96.2% vs 47.5%), Llama 3.3 70B (67.8% vs 26.9%), Claude 3.5 Sonnet (28.1% vs 3.3%). The same pattern shows up across all four domains—healthcare, finance, legal, corporate—which pins inter-agent communication as the weak link in today’s multi-agent implementations.
@@
To quantify the practical impact of internal channel vulnerabilities, we tested the hidden leakage hypothesis (H1), described earlier: runs exist where the final output (C1) is safe but internal channels (C2 or C5) leak sensitive data (Fig. 5). Formally, $H1=\neg C1_{leak}\land(C2_{leak}\lor C5_{leak})$—i.e., traces where output-only audits would pass but internal channels leaked.
Figure 5: Output-only audit gap (n=4,979 traces). 41.7% of violations occur in internal channels while final output passes checks—audits inspecting only C1 miss nearly half of all violations.
@@
TABLE IX: H1 Validation: Output-Only Audit Insufficiency (n=4,979). 95% Wilson score CIs. H1 = C1 safe but C2 or C5 leaked. The pattern holds consistently across all four domains.
@@
Table IX shows the H1 validation results (visualized in Fig. 6). Nearly half of all violations occur in internal channels while the final output passes checks, confirming that output-only audits are insufficient for multi-agent systems.
Figure 6: Multi-agent privacy violations breakdown (n=4,979 traces). Missed violations by output-only audit (H1 rate: 41.7%) across verticals. Higher-complexity domains (Finance, Legal) show elevated miss rates on internal channels.
@@
Table VI aggregates results across all 4,979 traces (all threat levels and defense configurations), whereas Table VII isolates A0 (benign) scenarios from the defense-evaluation subset exclusively. In the aggregate data, channel rates by threat level are: A0 (n=2,485n=2{,}485): C1=26.6%, C2=69.3%, C5=46.4%; A1 (n=1,234n=1{,}234): C1=28.4%, C2=70.6%, C5=48.2%; A2 (n=1,230n=1{,}230): C1=26.8%, C2=66.3%, C5=45.9%. The pattern C2>>C1 holds across all three threat levels. The higher C1 rate in Table VII (Vanilla: 48%) compared to the aggregate A0 rate (26.6%) arises because the defense-evaluation subset uses a smaller, curated set of A0 scenarios specifically designed to stress-test defense mechanisms and includes more disclosure-intensive tasks.
Table VII presents channel-specific effectiveness measurements on this defense-evaluation subset. Under baseline conditions without defenses, we observe 48% leakage on C1 and 31% on internal channels. Privacy-focused prompting reduces C1 to 19% (representing 60% effectiveness) while achieving only 6% reduction on internal channels. Output sanitization mechanisms demonstrate 98% effectiveness on C1 (reducing leakage to 1%) yet exhibit zero impact on internal channels, as these systems observe only final outputs rather than inter-agent communication or memory operations.
@@
[…]
@@
Table X summarizes framework-level vulnerabilities with no defenses enabled. All four frameworks CrewAI, AutoGPT, LangChain, MetaGPT—show similar C2 leakage (28–35%) without adversarial attacks. That baseline jumps to 68.8% once we add adversarial scenarios (A1/A2). The tight spread (28–35%) across architectures that differ radically CrewAI’s role-based delegation, AutoGPT’s recursive decomposition, LangChain’s chain abstractions, MetaGPT’s SOP workflows confirms the problem is systemic. None of these frameworks treat inter-agent privacy as a design goal.
@@
TABLE XI: Adversarial Attack Success Rates by Family.
@@
Table XI ranks attack families by success rate. Multi-agent coordination attacks (F4) achieve the highest success rate at 82.9%, exploiting inter-agent trust to use the coordination channel as an exfiltration vector and shared memory as a staging area for data extraction. Prompt/instruction attacks (F1) and tool-surface attacks (F2) are not far behind at 79.8% and 78.9% all comfortably above 75%.
@@
Log Leakage (C6) and the “Thought Gap”: System log leakage (C6) consistently runs higher than tool invocation leakage (C3). This reveals something interesting about agent behavior: they process sensitive information in their chain-of-thought (which ends up in logs) even when they ultimately decide not to pass that data to external tools. Claude 3.5 Sonnet shows the starkest example of this gap—11 percentage points between C3 (74%) and C6 (85%)—proving that even heavily safety-trained models can compromise privacy through their internal reasoning process.
Shadow Leakage: In over 65% of scenarios across all models, sensitive data escapes through C3 or C6 while C1 (user-facing output) stays clean. This is precisely the false sense of security that output-focused auditing creates—confidential information leaks through infrastructure channels that nobody is watching.
@@
We built a prototype interceptor that applies C1 redaction logic to internal channels C2 and C5. The test: 120 scenarios (30 per domain), stratified across A0/A1/A2 (40 each), run on CrewAI and LangChain with identical seeds and $\text{temperature}=0.7$. The interceptor masked vault fields outside the channel-specific allowed set and blocked raw vault transmission in inter-agent messages and memory writes.
Results: internal leakage dropped from 31.5% to 2.4% (95% CI: [0.4%, 6.8%]). External leakage stayed flat (8.2% vs 8.1%). Task Success Rate fell from 78.6% to 73.9%—a 4.7-point hit (95% CI: [1.2, 8.2]), mostly in scenarios needing legitimate data sharing between agents.
@@
Tier 1 (Canary Detection, 8%): Verbatim disclosure of our unique marker strings. Modern LLMs have largely learned to suppress such obvious markers—we found only 18 instances across the entire corpus.
Tier 2 (Structured Field Detection, 10%): Pattern matching for regulated identifiers—SSNs, credit card numbers, and the like. Safety training has helped here, bringing us to 24 instances. One caveat: this tier needs careful tuning because internal identifiers can trip SSN patterns (e.g., a ticket number like “TIC-987-65-4321” looks suspiciously like an SSN).
Tier 3 (Semantic Detection, 82%): The majority of detected violations fall into this category. In 184 cases, models rephrased sensitive information instead of quoting it verbatim—swapping “Patient is being treated for diabetes” for explicit ICD-10 codes, or saying “Account balance exceeds \$50,000” instead of the precise figure. Models do this because they are trying to be helpful, but helpfulness and confidentiality often pull in opposite directions.
@@
TABLE XII: Per-Model Channel Leak Rates (n=4,979 traces across five models).
@@
Figure 9: Per-channel leak rates (%) by domain. Important: These are channel-specific rates—the percentage of traces that leaked on each respective channel. A single trace can leak on multiple channels, so these rates do not sum to a scenario-level leakage rate. For example, Healthcare C2=87.1% means 87.1% of healthcare traces leaked via inter-agent messages, not that 87.1% of scenarios leaked overall. See Table IX for H1 (output-safe but internal-leaked) rates, which measure a different metric.
Note that Fig. 9 reports channel-specific rates (the percentage of traces leaking on each channel independently), while Table IX reports H1 rates (traces where C1 is safe but C2 or C5 leaked). These are complementary metrics measuring different aspects of leakage: the former quantifies per-channel vulnerability, the latter quantifies the audit gap.
As shown in Fig. 9, the domain analysis exhibits variation. Finance: lowest C1 rate at 19.8%—likely financial regulation-focused training data. Legal: highest C1 at 35.0%—higher disclosure rates in legal contexts. Healthcare shows intermediate rates (C1=24.7%) despite HIPAA and Law 25 concerns, while Corporate shows C1=29.3%. This variation across domains suggests domain-specific fine-tuning as a promising defense direction.
@@
Fig. 10 breaks this down by model and channel—internal channels (C2, C5) show higher leak rates than external output (C1) for all five models tested.
@@ VI-F Security-Utility Tradeoff Analysis
Analysis of 50 dedicated tradeoff scenarios shows TSR drops to 35–65% under strict privacy settings that prevent any sensitive data sharing between agents. Future defenses need selective disclosure—sharing only minimum necessary fields, enforced at the framework level with explicit data flow policies.
@@ VII-A Clarifying Internal Channel Leakage: Violations, Risks, and Breaches
(1) Policy violation (data minimization breach): An agent shares more data than the task strictly requires, violating the minimization principle (GDPR Art. 5(1)(c), Law 25 Sec. 5). For instance, a scheduler agent includes full medical history when only the appointment date matters. This is what our benchmark flags—the policy is broken even if no outsider ever sees the data.
@@
(3) Actual confidentiality breach: Sensitive data reaches an unauthorized party—an external attacker, an unintended third-party API, a user who should not have access. This is the classic definition of a breach.
@@
Tradeoff: This creates a difficult business decision. Our prototype shows that strict internal sanitization can drop Task Success Rate (TSR) by 4.7 points (from 78.6% to 73.9%). For highly regulated industries (e.g., healthcare), a $\sim$5% drop in utility is likely an acceptable cost for compliance assurance. However, for general-purpose consumer agents, this friction might be commercially prohibitive, creating a market bifurcation between ”secure enterprise agents” and ”unconstrained consumer agents.”
Implication: Organizations must decide their risk tolerance. If internal agents are fully trusted and logs are secured, (1) may be acceptable. If any agent could be compromised, or logs may be accessed by third parties, (1) becomes a precursor to (3). Our results inform this risk assessment; they do not prescribe a universal policy.
@@ VII-B Implications for System Designers
This work identifies architectural risks with regulatory implications; it does not constitute legal guidance. Multiple regulatory frameworks may be implicated. Under a strict reading of HIPAA’s minimum necessary standard, inter-agent messages containing protected health information could constitute disclosures that may require logging and access control. Quebec Law 25 mandates privacy-by-default (Section 9.1) and data minimization (Section 5)—our 57.8% internal leakage rate suggests potential non-compliance under this interpretation. GDPR Article 25 (Data Protection by Design) requires architectural safeguards that current frameworks lack; Articles 13–14 could be implicated when data subjects are unaware their information crosses agent boundaries. Under the California Consumer Privacy Act/California Privacy Rights Act (CCPA/CPRA) and PCI-DSS, inter-agent data flows may constitute “sharing” requiring disclosure, and persistent sensitive data in agent memory could violate storage requirements. The precise legal characterization of inter-agent data flows remains an open question for regulators and legal scholars.
@@ VII-C SDK Integration: Real-World CrewAI Showcase
To validate practical applicability, we built a fully-integrated showcase using the AgentLeak SDK with CrewAI. This experiment shows that hybrid detection (Presidio NER + LLM Judge) catches leakage in production-grade conditions—real task delegation, real tool usage.
Application Architecture and Tests. We built a portfolio management system with CrewAI v0.83+: three agents (Senior Research Analyst, Senior Financial Analyst, Investment Advisor). The integration tests process sensitive client data—IBAN, tax brackets, trade history—while agents must respect privacy constraints. The integration tests process real tool calls to market data APIs and internal calculators, providing a high-fidelity environment for privacy audit.
@@
Benchmark Configuration. Five LLMs via OpenRouter API: GPT-4o, GPT-4o-mini, Mistral Large 2411, Llama 3.3 70B Instruct, and Claude 3.5 Sonnet. Same workflow, same client data. Detection in HYBRID mode—Presidio patterns plus LLM Judge semantics.
Results. Table XIII presents results from the multi-model benchmark.
TABLE XIII: SDK Integration Benchmark: Privacy Leakage Across LLM Families (Single Complex Execution per model). C1=Final output, C2=Inter-agent, C3=Tool inputs, C4=Tool outputs, C5=Memory. Counts represent distinct leaked items per run. C1=0 confirms output-only audits would miss these violations. Contrast with Table XII which reports rates across 4,979 traces.
@@
What the numbers show. Every model leaked PII through internal channels (C2–C5)—no model family proved immune. All models sent sensitive data to external tools (C3) when the prompt suggested it; the privacy failure was baked into the workflow, not the model. Our hybrid detection caught everything: pattern-matched identifiers and semantically-detected vault fields alike. The consistency across model families confirms the vulnerability is architectural, not model-specific.

### v2

@@ VI-E Results
Each of the 4,979 traces contains both a single-agent execution (C1 only) and a multi-agent execution (C1+C2+C5) of the same scenario on the same model. This paired design enables direct comparison. Proportion of traces in which at least one sensitive field leaked through a given channel is used in Tables VII, IX, XIII. For high-level comparison, we use the arithmetic mean of C2 and C5 (57.8%) as a representative internal rate. The combined OR rate ($C2\lor C5$) is 68.8%.
@@
TABLE VIII: Defense Effectiveness by Channel Type (A0 Benign Scenarios Only). This table shows A0 (benign, no attack) scenarios only, where C1 leakage is higher due to models completing tasks without adversarial pressure. Under adversarial conditions (A1/A2), internal channels dominate as shown in Table VII. $\text{C1 Eff.}=(\text{Vanilla C1 Rate}-\text{Defense C1 Rate})/\text{Vanilla C1 Rate}\times 100\%$
@@
TABLE IX: Architecture Comparison: Single-Agent vs Multi-Agent (n=4,979 paired traces). Statistical test: McNemar’s test for paired proportions. C2-only = traces where C1 safe but C2 leaked; C5-only = traces where C1&C2 safe but C5 leaked. C5-only refers to the incremental leak rate; the +0.0% indicates that in our dataset, every memory leak (46.7%) co-occurred with an inter-agent (C2) or output (C1) leak
@@
Why the increase? Multi-agent systems create data flow pathways that did not exist before. Task delegation, shared context, result aggregation: all become channels where sensitive data can escape. The pattern is monotonic: add an agent, add a pathway. Single-agent systems have none of these; multi-agent systems turn them into potential leak points.
@@
What stands out is that C2 beats C1 for every single model we tested: GPT-4o (76.8% vs 17.2%), GPT-4o-mini (75.3% vs 41.2%), Mistral Large (96.2% vs 47.5%), Llama 3.3 70B (67.8% vs 26.9%), Claude 3.5 Sonnet (28.1% vs 3.3%). The same pattern shows up across all four domains (healthcare, finance, legal, corporate), which pins inter-agent communication as the weak link in today’s multi-agent implementations.
@@
To quantify the practical impact of internal channel vulnerabilities, we tested the hidden leakage hypothesis (H1), described earlier: runs exist where the final output (C1) is safe but internal channels (C2 or C5) leak sensitive data (Fig. 4). Formally, $H_{1}=\neg C1_{\mathrm{leak}}\land(C2_{\mathrm{leak}}\lor C5_{\mathrm{leak}})$, i.e., traces where output-only audits would pass but internal channels leaked.
Figure 4: Output-only audit gap (n=4,979 traces). 41.7% of violations occur in internal channels while final output passes checks; audits inspecting only C1 miss nearly half of all violations. Note: The 27.2% C1 leak rate and 41.7% H1 rate measure different metrics on the same population: C1 counts traces where the output leaked, while H1 counts traces where C1 was safe but internal channels (C2 or C5) leaked. These are not additive; they partition different subsets of the 4,979 traces.
@@
TABLE X: H1 Validation: Output-Only Audit Insufficiency (n=4,979). 95% Wilson score CIs. H1 = C1 safe but C2 or C5 leaked. The pattern holds consistently across all four domains
@@
Table X shows the H1 validation results (visualized in Fig. 5). Nearly half of all violations occur in internal channels while the final output passes checks, confirming that output-only audits are insufficient for multi-agent systems.
Figure 5: Multi-agent privacy violations breakdown (n=4,979 traces). Missed violations by output-only audit (H1 rate: 41.7%) across verticals. Higher-complexity domains (Finance, Legal) show elevated miss rates on internal channels. Note: these are independent H1 rates within each vertical and do not sum to 100%.
@@
Table VII aggregates results across all 4,979 traces (all threat levels and defense configurations), whereas Table VIII isolates A0 (benign) scenarios from the defense-evaluation subset exclusively. In the aggregate data, channel rates by threat level are: A0 (n=2,485n=2{,}485): C1=26.6%, C2=69.3%, C5=46.4%; A1 (n=1,234n=1{,}234): C1=28.4%, C2=70.6%, C5=48.2%; A2 (n=1,230n=1{,}230): C1=26.8%, C2=66.3%, C5=45.9%. The pattern C2>>C1 holds across all three threat levels. The higher C1 rate in Table VIII (Vanilla: 48%) compared to the aggregate A0 rate (26.6%) arises because the defense-evaluation subset uses a smaller, curated set of A0 scenarios specifically designed to stress-test defense mechanisms and includes more disclosure-intensive tasks.
Table VIII presents channel-specific effectiveness measurements on this defense-evaluation subset. Under baseline conditions without defenses, we observe 48% leakage on C1 and 31% on internal channels. Privacy-focused prompting reduces C1 to 19% (representing 60% effectiveness) while achieving only 6% reduction on internal channels. Output sanitization mechanisms demonstrate 98% effectiveness on C1 (reducing leakage to 1%) yet exhibit zero impact on internal channels, as these systems observe only final outputs rather than inter-agent communication or memory operations.
@@
TABLE XI: Framework-Level C2 Leakage (Representative Sample). Scope: A0 (benign) scenarios only, vanilla configuration, SDK integration tests. These rates (28–35%) represent the baseline without adversarial attacks. The 68.8% aggregate C2 rate in Table VII includes A1/A2 adversarial scenarios, explaining the ${\sim}2\times$ difference
@@
Table XI summarizes framework-level vulnerabilities with no defenses enabled. All four frameworks (CrewAI, AutoGPT, LangChain, MetaGPT) show similar C2 leakage (28–35%) without adversarial attacks. That baseline jumps to 68.8% once we add adversarial scenarios (A1/A2). The tight spread (28–35%) across architectures that differ radically (CrewAI’s role-based delegation, AutoGPT’s recursive decomposition, LangChain’s chain abstractions, MetaGPT’s SOP workflows) confirms the problem is systemic. None of these frameworks treat inter-agent privacy as a design goal.
@@
TABLE XII: Adversarial Attack Success Rates by Family
@@
Table XII ranks attack families by success rate. Multi-agent coordination attacks (F4) achieve the highest success rate at 82.9%, exploiting inter-agent trust to use the coordination channel as an exfiltration vector and shared memory as a staging area for data extraction. Prompt/instruction attacks (F1) and tool-surface attacks (F2) are not far behind at 79.8% and 78.9% all comfortably above 75%.
@@
Log Leakage (C6) and the “Thought Gap”: System log leakage (C6) consistently runs higher than tool invocation leakage (C3). This reveals something interesting about agent behavior: they process sensitive information in their chain-of-thought (which ends up in logs) even when they ultimately decide not to pass that data to external tools. Claude 3.5 Sonnet shows the starkest example of this gap, 11 percentage points between C3 (74%) and C6 (85%), proving that even heavily safety-trained models can compromise privacy through their internal reasoning process.
Shadow Leakage: In over 65% of scenarios across all models, sensitive data escapes through C3 or C6 while C1 (user-facing output) stays clean. This is precisely the false sense of security that output-focused auditing creates: confidential information leaks through infrastructure channels that nobody is watching.
@@
We built a prototype interceptor that applies C1 redaction logic to internal channels C2 and C5. The test: 120 scenarios (30 per domain), stratified across A0/A1/A2 (40 each), run on CrewAI and LangChain with identical seeds and temperature =0.7=0.7. The interceptor masked vault fields outside the channel-specific allowed set and blocked raw vault transmission in inter-agent messages and memory writes.
Results: internal leakage dropped from 31.5% to 2.4% (95% CI: [0.4%, 6.8%]). External leakage stayed flat (8.2% vs 8.1%). Task Success Rate fell from 78.6% to 73.9%, a 4.7-point hit (95% CI: [1.2, 8.2]), mostly in scenarios needing legitimate data sharing between agents.
@@
Tier 1 (Canary Detection, 8%): Verbatim disclosure of our unique marker strings. Modern LLMs have largely learned to suppress such obvious markers; we found only 18 instances across the entire corpus.
Tier 2 (Structured Field Detection, 10%): Pattern matching for regulated identifiers: SSNs, credit card numbers, and the like. Safety training has helped here, bringing us to 24 instances. One caveat: this tier needs careful tuning because internal identifiers can trip SSN patterns (e.g., a ticket number like “TIC-987-65-4321” looks suspiciously like an SSN).
Tier 3 (Semantic Detection, 82%): The majority of detected violations fall into this category. In 184 cases, models rephrased sensitive information instead of quoting it verbatim, swapping “Patient is being treated for diabetes” for explicit ICD-10 codes, or saying “Account balance exceeds \$50,000” instead of the precise figure. Models do this because they are trying to be helpful, but helpfulness and confidentiality often pull in opposite directions.
@@
TABLE XIII: Per-Model Channel Leak Rates (n=4,979 traces across five models)
@@
Figure 8: Per-channel leak rates (%) by domain. Important: These are channel-specific rates: the percentage of traces that leaked on each respective channel. A single trace can leak on multiple channels, so these rates do not sum to a scenario-level leakage rate. For example, Healthcare C2=87.1% means 87.1% of healthcare traces leaked via inter-agent messages, not that 87.1% of scenarios leaked overall. See Table X for H1 (output-safe but internal-leaked) rates, which measure a different metric.
Note that Fig. 8 reports channel-specific rates (the percentage of traces leaking on each channel independently), while Table X reports H1 rates (traces where C1 is safe but C2 or C5 leaked). These are complementary metrics measuring different aspects of leakage: the former quantifies per-channel vulnerability, the latter quantifies the audit gap.
As shown in Fig. 8, the domain analysis exhibits variation. Finance: lowest C1 rate at 19.8%, likely financial regulation-focused training data. Legal: highest C1 at 35.0%, higher disclosure rates in legal contexts. Healthcare shows intermediate rates (C1=24.7%) despite HIPAA and Law 25 concerns, while Corporate shows C1=29.3%. This variation across domains suggests domain-specific fine-tuning as a promising defense direction.
@@
Fig. 9 breaks this down by model and channel; internal channels (C2, C5) show higher leak rates than external output (C1) for all five models tested.
@@ VI-F Security-Utility Tradeoff Analysis
Analysis of 50 dedicated tradeoff scenarios shows TSR drops to 35–65% under strict privacy settings that prevent any sensitive data sharing between agents. Future defenses need selective disclosure, sharing only minimum necessary fields, enforced at the framework level with explicit data flow policies.
@@ VII-A Clarifying Internal Channel Leakage: Violations, Risks, and Breaches
(1) Policy violation (data minimization breach): An agent shares more data than the task strictly requires, violating the minimization principle (GDPR Art. 5(1)(c), Law 25 Sec. 5). For instance, a scheduler agent includes full medical history when only the appointment date matters. This is what our benchmark flags: the policy is broken even if no outsider ever sees the data.
@@
(3) Actual confidentiality breach: Sensitive data reaches an unauthorized party: an external attacker, an unintended third-party API, a user who should not have access. This is the classic definition of a breach.
@@
Tradeoff: This creates a difficult business decision. Our prototype shows that strict internal sanitization can drop Task Success Rate (TSR) by 4.7 points (from 78.6% to 73.9%). For highly regulated industries (e.g., healthcare), a $\sim$5% drop in utility is likely an acceptable cost for compliance assurance. However, for general-purpose consumer agents, this friction might be commercially prohibitive, creating a market bifurcation between “secure enterprise agents” and “unconstrained consumer agents.”
Implication: Organizations must decide their risk tolerance. If internal agents are fully trusted and logs are secured, (1) may be acceptable. If any agent could be compromised or logs may be accessed by third parties, (1) becomes a precursor to (3). Our results inform this risk assessment; they do not prescribe a universal policy.
@@ VII-B Implications for System Designers
This work identifies architectural risks with regulatory implications; it does not constitute legal guidance, and multiple regulatory frameworks may be implicated. Under a strict reading of HIPAA’s minimum necessary standard, inter-agent messages containing protected health information could constitute disclosures that may require logging and access control. Quebec Law 25 mandates privacy-by-default (Section 9.1) and data minimization (Section 5) – our 57.8% internal leakage rate suggests potential non-compliance under this interpretation. GDPR Article 25 (Data Protection by Design) requires architectural safeguards that current frameworks lack; Articles 13–14 could be implicated when data subjects are unaware their information crosses agent boundaries. Under the California Consumer Privacy Act/California Privacy Rights Act (CCPA/CPRA) and PCI-DSS, inter-agent data flows may constitute “sharing” requiring disclosure, and persistent sensitive data in agent memory could violate storage requirements. The precise legal characterization of inter-agent data flows remains an open question for regulators and legal scholars.
@@ VII-C SDK Integration: Real-World CrewAI Showcase
To validate practical applicability, we built a fully-integrated showcase using the AgentLeak SDK with CrewAI. This experiment shows that hybrid detection (Presidio Named Entity Recognition (NER) + LLM Judge) catches leakage in production-grade conditions: real task delegation, real tool usage.
Application Architecture and Tests. We built a portfolio management system with CrewAI v0.83+: three agents (Senior Research Analyst, Senior Financial Analyst, Investment Advisor). The integration tests process sensitive client data (IBAN, tax brackets, trade history) while agents must respect privacy constraints. The integration tests process real tool calls to market data APIs and internal calculators, providing a high-fidelity environment for privacy audit.
@@
Benchmark Configuration. Five LLMs via OpenRouter API: GPT-4o, GPT-4o-mini, Mistral Large 2411, Llama 3.3 70B Instruct, and Claude 3.5 Sonnet. Same workflow, same client data. Detection in HYBRID mode: Presidio patterns plus LLM Judge semantics.
Results. Table XIV presents results from the multi-model benchmark.
TABLE XIV: SDK Integration Benchmark: Privacy Leakage Across LLM Families (Single Complex Execution per model). Counts represent distinct leaked items per run. C1=0 confirms output-only audits would miss these violations. Contrast with Table XIII which reports rates across 4,979 traces
@@
What the numbers show. Every model leaked PII through internal channels (C2–C5); no model family proved immune. All models sent sensitive data to external tools (C3) when the prompt suggested it; the privacy failure was baked into the workflow, not the model. Our hybrid detection caught everything: pattern-matched identifiers and semantically-detected vault fields alike. The consistency across model families confirms the vulnerability is architectural, not model-specific.